OpenStudy (anonymous):

integrate cos(pi/x)/x^2 from [2,infinity]

OpenStudy (anonymous):

the answer is 1/pi im just not sure on the process

OpenStudy (dominusscholae):

\[\cos(\pi/x)/x^2\] Use u substitution with u being pi/x so that du = \[-\pi/x^2 dx\].

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

ok so i substituted in u and du but my solution was 1 where -(sin(pi/infinity)-sin(pi/2)

OpenStudy (anonymous):

*is equal to one

OpenStudy (dumbcow):

you forgot about pi \[du = -\pi/x^{2} dx\] so \[dx = -x^{2}/\pi du\] \[\rightarrow -1/\pi \int\limits_{2}^{\infty}\cos(u) du\] after integrating, you get 1/pi

OpenStudy (anonymous):

thanks so much!